OpenStudy (anonymous):

determine the 100th term of arithmetic sequence with t7=465 and t13=219. plz show me the procedure, thats what i have trouble with.

OpenStudy (eric_d):

a = the value of the first term, d = the difference between two consecutive terms Let's look at the 7th term:

OpenStudy (eric_d):

t7= a+6d=465...............1

OpenStudy (eric_d):

Then t13= a+12d=219................2

OpenStudy (eric_d):

Solve equations (1) and (2) to find a and d

OpenStudy (anonymous):

ahh, i see i can do the rest, thank you!

OpenStudy (eric_d):

welcome